# Mapping surface height dynamics to subsurface flow physics in free-surface turbulent flow using a shallow recurrent decoder

### v1

@@ Abstract
Near-surface turbulent flows beneath a free surface are reconstructed from sparse measurements of the surface height variation, by a novel neural network algorithm known as the SHallow REcurrent Decoder (SHRED). The reconstruction of turbulent flow fields from limited, partial, or indirect measurements remains a grand challenge in science and engineering. The central goal in such applications is to leverage easy-to-measure proxy variables in order to estimate quantities which have not been, and perhaps cannot in practice be, measured. Specifically, in the application considered here, the aim is to use a sparse number of surface height point measurements of a flow field, or drone video footage of surface features, in order to infer the turbulent flow field beneath the surface. SHRED is a deep learning architecture that learns a delay-coordinate embedding from a few surface height (point) sensors and maps it, via a shallow decoder trained in a compressed basis, to full subsurface fields, enabling fast, robust training from minimal data. We demonstrate the SHRED sensing architecture on both fully resolved DNS data and PIV laboratory data from a turbulent water tank. SHRED is capable of robustly mapping surface height fluctuations to full-state flow fields up to about two integral length scales deep, with as few as three surface measurements.
@@ I Introduction
A particular application for which data-driven methods have yielded recent progress is the detailed reconstruction of turbulent flows near interfaces, where measurements at the interface give input to the learning algorithm. Such cases include reconstructing the bottom topography [22] and full subsurface velocity fields [18] from free surface measurements, and reconstructing wall-bounded turbulence from wall measurements only. The latter using CNNs for large-scale reconstruction [23, 24], CNN-based variational autoencoders [20], or GANs [25, 19], yielding great reconstruction performance compared to linear methods. For the reconstruction of subsurface flow fields from free surface data, Xuan & Shen (2023) [18] applied a CNN on data from DNS, using surface elevation and the surface velocity field as input. The results are promising and outperform previous linear reconstruction models, indicating that turbulent flow fields in, say, a river or lake, may indeed be inferred from the water surface only. In this case, however, the CNN is applied image-by-image for reconstruction, hence independent of temporal dynamics. Moreover, it is pointed out that CNN methods may underestimate the fluctuating amplitudes of large-scale structures. Beyond these issues, CNN-based approaches often prove to be computationally expensive and either require dense surface fields or struggle with noisy experimental data [11], leaving a gap for a lightweight model that can operate on sparse field data and generalize across simulation and laboratory flows. A more general issue with ”sensing” turbulent flow fields from interface measurements is the fall in accuracy with increased distance from the measured interface, due to less correlation of the flow and interface far from the latter. Hence, closing the gaps of both efficiency and accuracy is essential to take subsurface flow sensing to a level of practical utility, i.e., for practical remote sensing of river turbulence and subsequent gas‑exchange estimation.
@@
This paper presents a case study of SHRED deployed to flow data from turbulent free-surface flow, obtained from two separate sources: direct numerical simulations and experiments. In both cases, isotropic, homogeneous turbulence is created well below the surface whence it naturally diffuses towards the surface. Fig. 1 exhibits the SHRED algorithm in relation to our mapping of surface measurements to subsurface flows. The present flow case is well documented in the turbulence literature, both experimentally [34, 35, 36, 37, 38, 39, 40, 41, 42] and numerically [43, 44, 45, 3, 2], and can be considered fundamental to our understanding of free-surface turbulence dynamics. The relevance to a number of naturally occurring flows is immediate, such as the near-surface flow of a river or the upper layer of the ocean when the free-surface shear stress due to wind is negligible. We present the performance of SHRED on data from this flow case as a proof-of-concept of data-driven sensing of subsurface flows from surface measurements only, in a highly challenging turbulent environment characterized by large intermittent structures and absence of periodic features. In stress testing SHRED for different Reynolds numbers from DNS and laboratory experiments, we aim at connecting the SHRED performance to real non-ideal flows as typically found in real-life cases of rivers and oceans. Moreover, our results highlight the significant experimental gain of reconstruction of subsurface flows from surface-only measurements, as a step towards more accurate non-intrusive measurement techniques.
The paper is organized as follows: In § II we introduce the reconstruction model SHRED and in § III we provide details on the simulations and experiments from which our training data and reconstruction data sets are gathered, as well as data compression and performance metrics used. We present our results in § IV, among them a set of error metrics which discloses how the performance of SHRED varies with distance from the reconstructed plane to the free surface. Lastly, we draw conclusions in § V and include an appendix with details on parameter tuning thereafter.
@@ III.1 DNS and experiments
The simulations mimic experiments of free-surface homogeneous turbulence where turbulence is generated far beneath the water surface and diffuses towards the surface through self-interaction of turbulence vortices and viscous diffusion, with turbulence generated by an oscillating grid [see, e.g., 34, 36, 38, 40] or jets with zero net flow [e.g. 41, 42, 53, 52]. In the simulations, the non-dimensional, incompressible Navier–Stokes equations and continuity equation are solved on a three-dimensional domain, with periodic boundary conditions in horizontal directions, a free-slip boundary at the bottom, and a free surface boundary on the top. The latter is enforced by the dynamic and kinematic boundary conditions and tracked by a surface-adhering grid that undulates in the vertical direction as the surface elevation varies. Turbulence is generated in the center region of the domain by random linear forcing [54, 55], modulated by a depth-dependent function which is strongest in the center of the domain and drops to zero well below the surface-influenced region (see Fig. 2a). The top region, denoted the ‘free region’, is free from random forcing. Only data from the free region is used in the analysis. Surface tension is neglected in the simulations, which has little effect on the degree of correlation between surface and bulk as long as the flow is in the high-Weber-number regime (see extensive discussion in [2]).
@@
In the experiments, turbulence is generated by randomly actuated water jets below the water surface in a rectangular tank (see Fig. 2(b); note that only the two bottom rows of the jets are active, see [52]). By a novel data collection technique that combines surface profilometry with particle image velocimetry (PIV), the surface displacement and velocity in a subsurface horizontal plane are captured simultaneously (see [51]; for alternative methods of measuring the surface and the subsurface simultaneously, see [56, 57, 58, 22, 59, 53]]). Unlike in the simulations, only a single velocity plane is resolved at each time step. The laser sheet in use is varied between four different levels between ensembles to get estimates for depth dependency. During characterization of the experiment, a vertical laser sheet is used and, hence, well resolved root-mean-square measures of the velocity are obtained. The experiments measured surface profilometry with each corresponding PIV velocity field at 20 intervals, or ensemble cases, each with a 1 minute duration. In order for the cases to be independent, there was a minute pause in between ensemble cases. Due to noise and issues in the profilometry for a few cases, only ensemble cases are considered usable data for the SHRED analysis.
Case details for the four data sets are listed in Table 1. There, two additional dimensionless numbers are included alongside the Reynolds number. These are the Forude number and the Weber number. While the Reynolds number expresses the ratio of inertial to viscous forces, the Froude number takes into account the gravitational acceleration, , and the Weber number expresses the influence of inertial forces to forces due to surface tension, . Take note that directly comparing DNS data and data from experiments is not trivial (see extensive discussion in [52]). With non-dimensional DNS data the closest we get to a direct comparison is the turbulent Reynolds number and the Taylor Reynolds number. Yet even here, alternatives in how to compute length scales introduce discrepancies. Consider, for example, how computing integral length scales from dissipation and characteristic velocity (used in the DNS data) or by structure functions (used in the experimental data) introduce a difference in the length scale, and thus also in the Reynolds number of turbulence, , where is the representative velocity (i.e., the root-mean-square velocity, see [60]), is the integral scale and is the kinematic viscosity. Details on how these quantities are calculated in practice can be found in refs. [2, 52].
@@ III.2 Compression
Figs. 3 and 4 show the SVD modes and turbulence power spectra for cases S2 (DNS) and E2 (experiment), respectively. Through SVD the flow data is decomposed in matrices that contain spatial (), temporal (), and energy () information (strictly speaking, the latter matrix contains the singular values, , which can be considered as a measure of energy for the flow patterns in and ). Decomposing a dataset of timesteps results in modes. For example, case S2 has modes; E2 has . The modes are ordered by energy/singular-value, hence . After decomposition, the exact flow data set can be reconstructed by matrix multiplication of and . Spatial and temporal coefficients of eight modes for a velocity field near the surface are shown in the upper and middle panels of the figs. 3 and 4 . We observe that the large-scale spatial modes have large singular values, meaning the majority of the energy content is found in these modes – as expected from our knowledge of the turbulent flow. Likewise, modes of higher order (smaller singular value/lower energy) are related to smaller structures, as seen from comparing, e.g., to in Fig. 3 . Moreover, we observe a more rapid variation in the temporal coefficients as the order is increased. This is as expected for small compared to large scales in the turbulence, yet it may also be a result of noise, which typically shows up in small-scale data, another argument for using a compressed rather than a full data set with SHRED.
@@
The choice of rank (i.e., the range of modes to retain) for the low-rank representation of the data is done on the basis of 1) the observance of the spatial and temporal modes and their scales and variance, 2) the rapid decline of singular values per rank number, 3) from the turbulence energy spectra dependency on the rank number, and 4) post-result evaluation of optimal values for best SHRED performance. From a computational perspective (points 1 & 2), the optimal low-rank representation contains the relevant modes and covers most of the cumulative sum of the singular values. From a fluid mechanics perspective (point 3), we want to choose a rank truncation that matches the original spectrum well into the upper part of the inertial range. The results in Figs. 3 and 4 suggest that rank 250 is sufficient for the S2 case, and rank 100 is sufficient for E2. This corresponds to a cutoff at normalized wavenumber . Hence, all scales down to a sixth of the integral length scale are resolved. A parametric study of SHRED performance for different rank truncations is included in the appendix A. We find that the performance of SHRED decreases past a certain rank value, with the exact value depending on the flow case. Taking this into account, we choose the low-rank representation of the data to be truncated as listed in Table 2. We note that for experimental cases, we choose a rather low rank, with only a cumulative sum of singular values of to percent. This is because the noise present in the experimental data makes it significantly harder to handle higher-rank data in these cases than in the data from the DNS.
| Case | Low-rank | Full-rank | Cumulative sum | Rank truncation | |
@@ III.3 Training, validation and testing
We deploy SHRED on datasets consisting of simultaneous velocities and free-surface elevation resolved in space and time, from the two DNS cases and two experimental cases according to the principles laid out in Fig. 1. We input time series of surface elevation from three randomly placed surface sensor points into a two-layer LSTM. The LSTM encodes these input sequences into a latent representation of their temporal dynamics. This latent vector is then passed to a shallow decoder network (SDN), which maps it onto the velocity fields across depth. We do this in compressed space, by feeding into the SDN the compressed matrices of the SVD decomposition for the surface elevation and the subsurface velocity fields. These fields are used in training and validation not to learn the subsurface time dynamics but only to learn the mapping of the surface time dynamics onto the subsurface fields. This is an essential detail in the context of remote sensing. To output the full spatiotemporal fields, we save the compressed and matrices from the training data, and matrix multiplication with then yields the reconstruction fields. This is computationally cheaper than feeding in the full spatiotemporal fields, although that is also possible.
Each SHRED model is independently trained on one continuous dataset at a time, that is, no cross-case or cross-domain training is used. For each case, we randomly split the data into 80% training, 10% validation, and 10% testing snapshots, all within the same time series. An ensemble of SHRED models is trained for each case to assess convergence and uncertainty for each run.
@@
Fig. 5 shows a typical validation MSE loss curve for cases S2 and E2. We observe a steep drop in validation after around epochs for case S2, and around epochs for case E2, and a convergence typically occurs after a few hundred epochs. Running on GPU, a full SHRED run of surface paired with a single velocity plane, takes 1-2 minutes on a regular desktop computer. A simultaneous run with all planes (only possible with the datasets in cases S1 and S2) takes up to 10 minutes.
@@ III.4.3 Power spectral density error
i.e. spectra are normalized by the maximum of the ground truth spectrum for that case. The PSD provides a measure of how energy is distributed across spatial length scales.
@@
We choose to exclude scales that are not reliably represented by the low-rank data used in training. Operationally, is set near the onset where the rank-truncated SVD spectrum begins to depart noticeably () from the ground-truth spectrum; in practice this falls in the intermediate (inertial-range) wavenumbers for all datasets. The same is used for ground truth and reconstructions within a given case. By setting these integration limits, rather than using the full spectrum, we neglect the major contribution in the error (as compared to the full-rank ground truth) the low-rank SVD truncation process itself would produce. Therefore, it is only relevant to compare the spectral power of the SHRED reconstruction and full rank ground truth down to the low-rank SVD-resolved spatial scales. For the experimental data, this corresponds to a wavelength resolution of cm, which is about 25% of the integral length scales. For the DNS data, the cutoff wavelength is chosen at around 17% of the integral length scale.

### v2

@@ Abstract
Near-surface turbulent flows beneath a free surface are reconstructed from sparse measurements of the surface height variation, by a novel neural network algorithm known as the SHallow REcurrent Decoder (SHRED). The reconstruction of turbulent flow fields from limited, partial, or indirect measurements remains a grand challenge in science and engineering. The central goal in such applications is to leverage easy-to-measure proxy variables in order to estimate quantities which have not been, and perhaps cannot in practice be, measured. Specifically, in the application considered here, the aim is to use a sparse number of surface height point measurements of a flow field, or drone video footage of surface features, in order to infer the turbulent flow field beneath the surface. SHRED is a deep learning architecture that learns a delay-coordinate embedding from a few surface height (point) sensors and maps it, via a shallow decoder trained in a compressed basis, to full subsurface fields, enabling fast, robust training from minimal data. We demonstrate the SHRED sensing architecture ontwo types of turbulent data from recent studies (Aarnes et al. J. Fluid Mech. 1007 A38, 2025 and Babiker et al. arXiv:251003732, 2025, respectively): fully resolved DNS data and PIV laboratory data from a turbulent water tank. SHRED is capable of robustly mapping surface height fluctuations to full-state flow fields up to about two integral length scales deep, with as few as three surface measurements.
@@ I Introduction
A particular application for which data-driven methods have yielded recent progress is the detailed reconstruction of turbulent flows near interfaces, where measurements at the interface give input to the learning algorithm. Such cases include reconstructing the bottom topography [23] and full subsurface velocity fields [64] from free surface measurements, and reconstructing wall-bounded turbulence from wall measurements only. The latter using CNNs for large-scale reconstruction [26, 25], CNN-based variational autoencoders [34], or GANs [25, 13], yielding great reconstruction performance compared to linear methods. For the reconstruction of subsurface flow fields from free surface data, Xuan & Shen (2023) [64] applied a CNN on data from DNS, using surface elevation and the surface velocity field as input. The results are promising and outperform previous linear reconstruction models, indicating that turbulent flow fields in, say, a river or lake, may indeed be inferred from the water surface only. In this case, however, the CNN is applied image-by-image for reconstruction, hence independent of temporal dynamics. Moreover, it is pointed out that CNN methods may underestimate the fluctuating amplitudes of large-scale structures.
Beyond these issues, CNN-based approaches often prove to be computationally expensive and either require dense surface fields or struggle with noisy experimental data [17], leaving a gap for a lightweight model that can operate on sparse field data and generalize across simulation and laboratory flows. A more general issue with ”sensing” turbulent flow fields from interface measurements is the fall in accuracy with increased distance from the measured interface, due to less correlation of the flow and interface far from the latter. Hence, closing the gaps of both efficiency and accuracy is essential to take subsurface flow sensing to a level of practical utility, i.e., for practical remote sensing of river turbulence and subsequent gas‑exchange estimation.
@@
This paper presents a case study of SHRED deployed to flow data from turbulent free-surface flow, made available to us from two separate sources: direct numerical simulations performed at the University of Minnesota, previously published in [1, 2] and experiments from a recent laboratory study at the Norwegian University of Science and technology [3, 4]. In both datasets isotropic, homogeneous turbulence was created well below the surface whence it naturally diffuses towards the surface. Fig. 1 exhibits the SHRED algorithm in relation to our mapping of surface measurements to subsurface flows. This category of free-surface turbulent flow is well documented in the turbulence literature, both experimentally [54, 33, 8, 41, 30, 12, 39, 56, 36] and numerically [28, 31, 32, 5, 1], and can be considered fundamental to our understanding of free-surface turbulence dynamics. The relevance to a number of naturally occurring flows is immediate, such as the near-surface flow of a river or the upper layer of the ocean when the free-surface shear stress due to wind is negligible. We present the performance of SHRED on such free-surface turbulent flow data as a proof-of-concept of data-driven sensing of subsurface flows from surface measurements only, in a highly challenging turbulent environment characterized by large intermittent structures and absence of periodic features. In stress testing SHRED for different Reynolds numbers from DNS and laboratory experiments, we aim at connecting the SHRED performance to real non-ideal flows as typically found in natural environments such as rivers and oceans. Moreover, our results highlight the significant experimental gain of reconstruction of subsurface flows from surface-only measurements, as a step towards more accurate non-intrusive measurement techniques.
The paper is organized as follows: In § II we introduce the reconstruction model SHRED and in § III we summarize the simulations and experiments from which our training data and reconstruction datasets originate, as well as data compression and performance metrics used. We present our results in § IV, among them a set of error metrics which discloses how the performance of SHRED varies with distance from the reconstructed plane to the free surface. Lastly, we draw conclusions in § V and include an appendix with details on parameter tuning thereafter.
@@ III.1 DNS and experiments
The simulations mimic experiments of free-surface homogeneous turbulence where turbulence is generated far beneath the water surface and diffuses towards the surface through self-interaction of turbulence vortices and viscous diffusion, with turbulence generated by an oscillating grid [see, e.g., 54, 8, 30, 39] or jets with zero net flow [e.g. 56, 36, 48, 3]. In the simulations [1, 61, 28], the non-dimensional, incompressible Navier–Stokes equations and continuity equation were solved on a three-dimensional domain, with periodic boundary conditions in horizontal directions, a free-slip boundary at the bottom, and a free surface boundary on the top. The latter is enforced by the dynamic and kinematic boundary conditions and tracked by a surface-adhering grid that undulates in the vertical direction as the surface elevation varies. Turbulence is generated in the center region of the domain by random linear forcing [40, 47], modulated by a depth-dependent function which is strongest in the center of the domain and drops to zero well below the surface-influenced region (see Fig. 2a). The top region, denoted the ‘free region’, is free from random forcing. Only data from the free region is used in the analysis. Surface tension is neglected in the simulations, which has little effect on the degree of correlation between surface and bulk as long as the flow is in the high-Weber-number regime (see extensive discussion in [1]).
@@
The experimental flow data is from the experiment of Babiker et al. [3] where full details can be found. Turbulence was generated by randomly actuated water jets below the water surface in a rectangular tank (see Fig. 2(b); note that only the two bottom rows of the jets were active.). By a data collection technique developed recently by Semati et al [49] that combines surface profilometry with particle image velocimetry (PIV), the surface displacement and velocity in a subsurface horizontal plane were captured simultaneously. Unlike in the simulations, only a single velocity plane was resolved at a time, serially measuring at four different depth levels between ensembles to get estimates for depth dependency. During characterization of the experiment, a vertical laser sheet was used and, hence, well resolved root-mean-square measures of the velocity are obtained. The experimental dataset consists of the measured free-surface elevation directly above each measured PIV velocity field for 20 time intervals intervals, or ensemble cases, each with a 1-minute duration. Of these, were considered suitable data for the SHRED analysis we perform here.
Flow metrics for the four datasets are listed in Table 1. There, two additional dimensionless numbers are included alongside the Reynolds number. These are the Froude number and the Weber number. While the Reynolds number expresses the ratio of inertial to viscous forces, the Froude number takes into account the gravitational acceleration, , and the Weber number expresses the influence of inertial forces to forces due to surface tension, . Take note that directly comparing DNS data and data from experiments is not trivial (see extensive discussion in [3]). With non-dimensional DNS data the closest we get to a direct comparison is the turbulent Reynolds number and the Taylor Reynolds number. Yet even here, alternatives in how to compute length scales introduce discrepancies. Consider, for example, how computing integral length scales from dissipation and characteristic velocity (used in the DNS data) or by structure functions (used in the experimental data) introduce a difference in the length scale, and thus also in the Reynolds number of turbulence, , where is the representative velocity (i.e., the root-mean-square velocity, see [53]), is the integral scale and is the kinematic viscosity. Details on how these quantities are calculated in practice can be found in refs. [1, 3].
@@ III.2 Compression
Figs. 3 and 4 show the SVD modes and turbulence power spectra for datasets S2 (DNS) and E2 (experiment), respectively. Through SVD the flow data is decomposed in matrices that contain spatial (), temporal (), and energy () information (strictly speaking, the latter matrix contains the singular values, , which can be considered as a measure of energy for the flow patterns in and ). Decomposing a dataset of timesteps results in modes. For example, data case S2 has modes; E2 has . The modes are ordered by energy/singular-value, hence . After decomposition, the exact flow dataset can be reconstructed by matrix multiplication of and . Spatial and temporal coefficients of eight modes for a velocity field near the surface are shown in the upper and middle panels of the figs. 3 and 4 . We observe that the large-scale spatial modes have large singular values, meaning the majority of the energy content is found in these modes—as expected from our knowledge of the turbulent flow. Likewise, modes of higher order (smaller singular value/lower energy) are related to smaller structures, as seen from comparing, e.g., to in Fig. 3 . Moreover, we observe a more rapid variation in the temporal coefficients as the order is increased. This is as expected for small compared to large scales in the turbulence, yet it may also be a result of noise, which typically shows up in small-scale data, another argument for using a compressed rather than a full dataset with SHRED.
@@
The choice of rank (i.e., the range of modes to retain) for the low-rank representation of the data is done on the basis of 1) the observance of the spatial and temporal modes and their scales and variance, 2) the rapid decline of singular values per rank number, 3) from the turbulence energy spectra dependency on the rank number, and 4) post-result evaluation of optimal values for best SHRED performance. From a computational perspective (points 1 & 2), the optimal low-rank representation contains the relevant modes and covers most of the cumulative sum of the singular values. From a fluid mechanics perspective (point 3), we want to choose a rank truncation that matches the original spectrum well into the upper part of the inertial range. The results in Figs. 3 and 4 suggest that rank 250 is sufficient for the S2 data case, and rank 100 is sufficient for E2. This corresponds to a cutoff at normalized wavenumber . Hence, all scales down to a sixth of the integral length scale are resolved. A parametric study of SHRED performance for different rank truncations is included in the appendix A. We find that the performance of SHRED decreases past a certain rank value, with the exact value different for each flow. Taking this into account, we choose the low-rank representation of the data to be truncated as listed in Table 2. We note that for experimental data cases, we choose a rather low rank, with only a cumulative sum of singular values of to percent. This is because the noise present in the experimental data makes it significantly harder to handle higher-rank data than in the data from the DNS.
| Dataset | Low-rank | Full-rank | Cumulative sum | Rank truncation | |
@@ III.3 Training, validation and testing
We deploy SHRED on the four datasets in Tab. 1 consisting of simultaneous velocities and free-surface elevation resolved in space and time, according to the principles laid out in Fig. 1. We input time series of surface elevation from three randomly placed surface sensor points into a two-layer LSTM. The LSTM encodes these input sequences into a latent representation of their temporal dynamics. This latent vector is then passed to a shallow decoder network (SDN), which maps it onto the velocity fields across depth. We do this in compressed space, by feeding into the SDN the compressed matrices of the SVD decomposition for the surface elevation and the subsurface velocity fields. These fields are used in training and validation not to learn the subsurface time dynamics but only to learn the mapping of the surface time dynamics onto the subsurface fields. This is an essential detail in the context of remote sensing. To output the full spatiotemporal fields, we save the compressed and matrices from the training data, and matrix multiplication with then yields the reconstruction fields. This is computationally cheaper than feeding in the full spatiotemporal fields, although that is also possible.
Each SHRED model is independently trained on one continuous dataset at a time, that is, no cross-case or cross-domain training is used. For each dataset, we randomly split the data into 80% training, 10% validation, and 10% testing snapshots, all within the same time series. An ensemble of SHRED models is trained in each case to assess convergence and uncertainty for each run.
@@
Fig. 5 shows a typical validation MSE loss curve for datasets S2 and E2. We observe a steep drop in validation after around epochs for case S2, and around epochs for E2, and a convergence typically occurs after a few hundred epochs. Running on GPU, a full SHRED run of surface paired with a single velocity plane, takes 1-2 minutes on a regular desktop computer. A simultaneous run with all planes (only possible with the datasets S1 and S2) takes up to 10 minutes.
@@ III.4.3 Power spectral density error
i.e. spectra are normalized by the maximum of the ground truth spectrum in each case. The PSD provides a measure of how energy is distributed across spatial length scales.
@@
We choose to exclude scales that are not reliably represented by the low-rank data used in training. Operationally, is set near the onset where the rank-truncated SVD spectrum begins to depart noticeably () from the ground-truth spectrum; in practice this falls in the intermediate (inertial-range) wavenumbers for all datasets. The same is used for ground truth and reconstructions within a given dataset. By setting these integration limits, rather than using the full spectrum, we neglect the major contribution in the error (as compared to the full-rank ground truth) the low-rank SVD truncation process itself would produce. Therefore, it is only relevant to compare the spectral power of the SHRED reconstruction and full rank ground truth down to the low-rank SVD-resolved spatial scales. For the experimental data, this corresponds to a wavelength resolution of cm, which is about 25% of the integral length scales. For the DNS data, the cutoff wavelength is chosen at around 17% of the integral length scale.